OpenStudy (rebeccapink6398):

WILL MEDAL+FAN!!! how do i cross multiply? PLEASE DO NOOOOT give me the answer. I would like to learn it :) 13/x = 39/87 (thirtyteen over 'X')(thirtynine over eightyseven)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

"cross multiply" as the math teachers say

OpenStudy (rebeccapink6398):

i dont know how to cross multiply, i forgot how :(

OpenStudy (anonymous):

\[\frac{a}{b}=\frac{c}{d}\\ a\times d = b\times c\]

OpenStudy (anonymous):

\[39\times x=87\times 13\]

OpenStudy (anonymous):

then \[x=\frac{87\times 13}{39}\]

OpenStudy (rebeccapink6398):

i dont understand that ?